Solve for the following systems of equations. If they cannot be solved, explain why.

5x-2y=-24
-4y=-48-10x

Answer by Lukas007(2)

5x-2y=-24
-4y=-48-10x
5x=-24+2y /*2
-4y=-48-10x
10x=-48+4y
-4y=-48+24-2y
10x=-48+4y
-2y=-24
10x=-48+4y
y=12
10x=-48+48
y=12
x=0
y=12